OpenStudy (anonymous):

find the exact value of: arctan(-square root of 3)

OpenStudy (anonymous):

you are looking for a number ( or angle if you prefer) whose tangent is \[-\sqrt{3}\]

OpenStudy (anonymous):

since a familiar triangle has one side \[\frac{\sqrt{3}}{2}\] and other side \[\frac{1}{2}\] and since \[\tan(x)=\frac{\sin(x)}{\cos(x)}\] this should be easy to find. make sure you are in the right quadrant